IONIZATION OF H_{2} BY X-RAYS IN THE CENTRAL MOLECULAR ZONE OF THE GALACTIC CENTER

\underline{MASAHIRO NOTANI} and TAKESHI OKA, Department of Astronomy and Astrophysics and Department of Chemistry, the Enrico Fermi Institute, the University of Chicago, Chicago, Illinois, 60637, USA.

Recent studies of the Galactic center using the infrared spectrum of H_{3}^{+} have revealed a high ionization rate of H_{2} on the order of ζ ∼ 3×10^{−15} s^{−1} in wide regions of the Central Molecular Zone (CMZ), a region with a radius of ∼ 150 pc at the Galactic center.^{a}^{b} So far we have ascribed this ionization rate, which is an order of magnitude higher than in the Galactic disk, to cosmic rays because of a high density of supernova remnants in the CMZ. In view of the abundant intense X-ray sources from 1 keV to several 10 keV in the region, however, there may be a significant ionization by X-rays also.

We estimate the ionization rate due to X-rays based on the large scale ART-P X-ray map of the Galactic center region.^{c} The calculations proceed in two steps. First we allow for the attenuation of the observed X-rays by the foreground gas to obtain the original intensities of the X-ray sources.^{d} We then use the corrected X-ray flux to calculate ionization rates of H_{2} in the CMZ. The calculation is also related to the heating of the gas by X-rays. Discussion of the details of calculations and the results will be presented.

aOka, T., Geballe, T. R., Goto, M., Usuda, T., and McCall, B. J. 2005, ApJ, 632 882

bGeballe, T. R., and Oka, T. 2010, ApJ, 709 L70.

cPavlinskii, M. N., Grebenev, S. A., and Syunyaev, R. A. 1992, Sov. Astron. Lett., 18 116.

dMorrison, R. and McCammon, D. 1983, ApJ, 270 119.